OpenStudy (anonymous):

can someone help me step by step with... the derivative of y=7cos^4(x)

OpenStudy (kropot72):

You can rewrite the equation as follows: \[y=7(\cos x)^{4}\] Then use the chain rule to find the derivative.

OpenStudy (babyslapmafro):

It might help to look at the problem in this format. \[y=7(cosx)^{4}\]

OpenStudy (kropot72):

@monroe17 Can you use the chain rule?

OpenStudy (anonymous):

28cos(x)*-sin(x)

OpenStudy (anonymous):

wait hold on i messed up

OpenStudy (anonymous):

what'd i do wrong?...

OpenStudy (anonymous):

nvm got it

OpenStudy (anonymous):

28cos^3(x)*-sin(x)

OpenStudy (anonymous):

nope its wrong ;/ what'd i do lol?

OpenStudy (kropot72):

\[\frac{dy}{dx}=\frac{d.7(\cos x)^{4}}{d.\cos x}\times \frac{d(\cos x)}{dx}\]

OpenStudy (kropot72):

@monroe17 Your last posting is correct.

OpenStudy (anonymous):

is 28cos^3(x)-sin(x) the same as -28cos^3(x)sin(x)?

OpenStudy (anonymous):

28cos^3(x)*-sin(x)

OpenStudy (sirm3d):

it is. you just forgot to multiply 28 and -1.

OpenStudy (anonymous):

oh okay thanks!

OpenStudy (sirm3d):

\[28(\cos x)^3(-sin x)\] should be clearer.